# A Unified Spectral Framework for Aging, Heterogeneous, and Distributed Order Systems via Weighted Weyl-Sonine Operators

### v1

@@ Abstract
Standard fractional calculus has successfully modeled systems with power-law memory. However, complex phenomena in heterogeneous media often exhibit multi-scale memory effects and aging properties that classical operators cannot capture. In this work, we construct a unified framework by defining the Weighted Weyl-Sonine Operators. This formalism offers a fundamental generalization of fractional calculus, freeing the theory from the constraints of power-law memory (via Sonine kernels), time-translation invariance (via scale and weight functions), and artificial history truncation (via Weyl integration).
The main result is a Generalized Spectral Mapping Theorem, proving that the Weighted Fourier Transform acts as a universal diagonalization map for these operators. We rigorously characterize the admissible memory kernels through the class of Complete Bernstein Functions, ensuring that the resulting operators preserve the fundamental properties of positivity and monotonicity. Furthermore, we establish a theoretical bridge between the algebraic Sonine definition and the analytical Marchaud representation involving Lévy measures.
Finally, we apply this theory to solve generalized relaxation equations and Weighted Distributed Order evolution problems, demonstrating that phenomena of ultra-slow diffusion and retarded aging can be treated explicitly within this unified spectral framework.
@@ 1 Introduction
To address these limitations, the theory of General Fractional Calculus (GFC) has emerged as a rigorous generalization, replacing the standard power-law with arbitrary kernels $k(t)$ satisfying the Sonine condition [2]. Yet, a fundamental question remains regarding the domain of these operators. As eloquently argued by Stinga [4], true “memory effects”—likened to an elephant never forgetting its complete history—are best captured by operators of the Weyl-Marchaud type defined on the entire real line $\mathbb{R}$. Restricting analysis to $t>0$ artificially truncates the system’s history, whereas a complete thermodynamic description requires integrating over $(-\infty,t)$.
Recently, significant strides have been made in understanding the structure of these generalized operators. A major breakthrough was achieved by Al-Refai and Fernandez [10], who provided a unified algebraic framework for Sonine operators via the method of conjugations. Their work elegantly demonstrates that many weighted operators can be constructed by conjugating standard derivatives with invertible linear operators. However, their approach implies an algebraic treatment typically using the Laplace transform on semi-axes. Consequently, the spectral analysis of “full history” operators on the unbounded domain $\mathbb{R}$—particularly those combining general memory kernels with time-dependent weights—remains an open problem. While abstract approaches to evolution equations have recently appeared (e.g., Lizama and Murillo-Arcila, 2025 [23]), an explicit harmonic analysis framework allowing for concrete calculations in these heterogeneous settings is still lacking.
This gap has led to a fragmentation of the field into a “zoo” of disconnected operators. Researchers must often choose between the power-law memory of Riemann-Liouville, the logarithmic scaling of Hadamard, the non-singular kernels of Caputo-Fabrizio, or tempered behaviors. While the Generalized Fractional Calculus of Luchko provides the algebraic structure for arbitrary kernels, the physical modeling of aging viscoelasticity, inertial oscillation, and ultra-slow diffusion demands a framework that simultaneously handles general memory geometry and time-domain heterogeneity.
@@ 3.1 Admissible Kernels and Operators
  • Case $n>1$ (High-Order): This allows for kernels that are continuous at the origin (regular kernels), as discussed in Remark LABEL:rem:singularity_discussion. In this setting, the operator $\mathfrak{D}_{\psi,\omega}^{(\kappa)}$ acts as a weighted derivative of order $n$, involving the $n$-th derivative of the convolution.
@@ 1. Explicit Form: The operator is defined pointwise by the weighted integral:
The construction in Definitions 3.6-3.7 hereditarily preserves the convolution algebra of the standard Sonine calculus, albeit non-trivially deformed by the diffeomorphism $\psi$ and the weight $\omega$.
@@
Table 1: Special cases of the Weighted Weyl-Sonine Operator $\mathfrak{D}_{\psi,\omega}^{(k)}$ acting on $\mathbb{R}$ (or valid sub-domains).
@@
Katugampola-type $\frac{z^{-\alpha}}{\Gamma(1-\alpha)}$ $t^{\rho}$ 11
@@
Prabhakar (Generalized) $z^{\beta-1}E_{\rho,\beta}^{\gamma}(-\lambda z^{\rho})$ $t$ 11
@@
Aging Viscoelasticity $\frac{z^{-\alpha}}{\Gamma(1-\alpha)}$ $\psi(t)$ $\omega(t)$
@@ 3.2 Spectral Analysis
Proposition 3.13 (Existence and Analytic Structure of Admissible Pairs).
Let $\Phi(s)$ be a function analytic in the right half-plane. A unique pair of kernels $(k,\kappa)$ satisfying the Sonine relation exists if and only if they are determined by the inverse Laplace transforms of (27):
@@
Moreover, to ensure physical realizability (non-negativity and passivity), if $\Phi(s)$ belongs to the class of Complete Bernstein Functions (CBF), then:
@@
    The derivative kernel $\kappa$ is a tempered distribution representing a Lévy measure (possibly with a drift), ensuring the operator models a sub-diffusive or relaxational process.
@@
    The integral kernel $k(t)$ is a locally integrable, completely monotone function on $(0,\infty)$, i.e.,
    $(-1)^{j}k^{(j)}(t)\geq 0$ for all $j\in\mathbb{N}_{0}$.
@@
Algebraic Existence. The convolution condition $(k*\kappa)(t)=t_{+}^{n-1}/(n-1)!$ is strictly equivalent in the Laplace domain to the algebraic equation $\hat{k}(s)\hat{\kappa}(s)=s^{-n}$. Identifying $\hat{\kappa}(s)=\Phi(s)$ immediately yields the unique solution for $\hat{k}(s)$ given in (26).
Analytic Properties. Assume $\Phi\in\mathcal{CBF}$. By the characterization of Complete Bernstein Functions (see Schilling et al. [20], Thm 6.2), the reciprocal $1/\Phi(s)$ is a Stieltjes function. Stieltjes functions $f(s)$ possess the integral representation:
$f(s)=\frac{a}{s}+b+\int_{0}^{\infty}\frac{1}{s+r}\,d\sigma(r),\quad a,b\geq 0,\,\sigma\geq 0.$
It is a classical result that for any Stieltjes function $f(s)$, the product $s^{-n}f(s)$ (for $n\geq 1$) remains the Laplace transform of a completely monotone function. Thus, the kernel $k(t)=\mathcal{L}^{-1}\{s^{-n}/\Phi(s)\}$ satisfies $(-1)^{j}k^{(j)}(t)\geq 0$.
This monotonicity ensures that the integral operator $\mathfrak{I}_{\psi,\omega}^{(k)}$ is positivity-preserving, a necessary condition for the validity of maximum principles and the well-posedness of the associated Cauchy problem. ∎
We now proceed to the main spectral mapping theorem. It is worth noting that the classical fractional calculus algebra—characterized by the power law $s^{-\alpha}$—is recovered here as a particular instance, now generalized by the spectral symbol $\Phi(s)$.
@@
Let $n\in\mathbb{N}$ and let $u\in\mathcal{S}_{\psi,\omega}(\mathbb{R})$. Let $(k,\kappa)$ be an admissible Sonine pair with associated Laplace symbol $\Phi(s):=\hat{\kappa}(s)$. The Weighted Weyl-Sonine operators satisfy the following spectral relations:
@@
Since $v\in\mathcal{S}(\mathbb{R})$, we apply the classical Convolution Theorem. Using the Sonine condition $\hat{k}(s)\hat{\kappa}(s)=s^{-n}$, we relate the symbol of $k$ to $\Phi$:
@@
The Generalized Spectral Mapping Theorem implies that the natural energy spaces for these operators are the Weighted Sonine-Sobolev spaces, defined via the spectral magnitude of the symbol $\Phi$:
$\mathcal{H}_{\psi,\omega}^{\Phi}(\mathbb{R}):=\left\{u\in L^{2}_{\psi,\omega}(\mathbb{R}):\int_{\mathbb{R}}\left(1+|\Phi(i\xi)|^{2}\right)|\mathcal{F}_{\psi,\omega}u(\xi)|^{2}\,d\xi<\infty\right\}.$
@@ 3.3 The Generalized Marchaud Representation
4 Applications to Anomalous Relaxation and Diffusion
In this final section, we demonstrate the versatility of the Weighted Weyl-Sonine framework by solving two classes of problems that arise in complex media: relaxation equations with oscillating memory and distributed-order evolution equations.
@@ 4 Applications to Anomalous Relaxation and Diffusion
While the theoretical framework developed in Section 3 supports operators of arbitrary order $n$, in this section we focus on the case $n=1$. This covers the vast majority of physical relaxation and diffusion phenomena, where the fundamental dynamics are governed by first-order evolution equations perturbed by memory.
4.1 Spectral Solvability of Generalized Relaxation Models
We consider the fundamental evolution equation governing the relaxation of a field $u\in\mathcal{S}_{\psi,\omega}$ (e.g., stress in a viscoelastic material or density in a porous medium) under the action of the generalized derivative.
@@ 4.1 Spectral Solvability of Generalized Relaxation Models
where $\lambda>0$ is a relaxation parameter, $f\in L^{2}_{\psi,\omega}(\mathbb{R})$ is a source term, and $\mathfrak{D}_{\psi,\omega}^{(\kappa)}$ is the Weighted Weyl-Sonine derivative (of order $n=1$) associated with the kernel $\kappa$.
@@
Let $\Phi(s)=\mathcal{L}\{\kappa\}(s)$ be the spectral symbol of the derivative. Assume the following Ellipticity Condition holds on the imaginary axis:
@@
Applying the Weighted Fourier Transform $\mathcal{F}_{\psi,\omega}$ to Eq. (32) and utilizing Theorem 3.14 (with $n=1$), the differential equation converts into an algebraic one in the frequency domain:
@@
Under condition (33), the multiplier $m(\xi)=(i\xi\Phi(i\xi)+\lambda)^{-1}$ belongs to $L^{\infty}(\mathbb{R})$. Since the Weighted Fourier Transform is a unitary isomorphism between $L^{2}_{\psi,\omega}$ and $L^{2}$, the solution is uniquely determined in the spectral domain. The explicit time-domain representation (34) follows directly from the inverse structure $\mathcal{F}_{\psi,\omega}^{-1}=M_{\omega}^{-1}Q_{\psi}\mathcal{F}^{-1}Q_{\psi}^{-1}M_{\omega}$, which maps the product of symbols back to the weighted convolution integral. ∎

### v2

@@ Abstract
While General Fractional Calculus has successfully expanded the scope of memory operators beyond power-laws, standard formulations remain predominantly restricted to the half-line via Riemann-Liouville or Caputo definitions. This constraint artificially truncates the system’s history, limiting the thermodynamic consistency required for modeling processes on unbounded domains. To overcome these barriers, we construct the Weighted Weyl-Sonine Framework, a generalized formalism that extends non-local theory to the entire real line without history truncation.
Unlike recent algebraic approaches based on conjugation for finite intervals, we develop a rigorous harmonic analysis framework. Our central contribution is the Generalized Spectral Mapping Theorem, which establishes the Weighted Fourier Transform as a unitary diagonalization map for these operators. This result allows us to rigorously classify and solve distinct physical regimes under a single algebraic structure. We explicitly derive exact solutions for diffusive relaxation (governed by Complete Bernstein Functions), inertial wave propagation (exhibiting oscillatory dynamics), and retarded aging (via distributed order), proving that our framework unifies the description of anomalous transport and wave mechanics in complex, time-deformed media.
@@ 1 Introduction
To address these limitations, the theory of General Fractional Calculus (GFC) has emerged as a rigorous generalization, replacing the standard power-law with arbitrary kernels $k(t)$ satisfying the Sonine condition [2]. Yet, a fundamental question remains regarding the domain of these operators. As eloquently argued by Stinga [4], true “memory effects”—likened to an elephant never forgetting its complete history—are best captured by operators of the Weyl-Marchaud type defined on the entire real line $\mathbb{R}$. Restricting analysis to $t>0$ artificially truncates the system’s history, whereas a complete thermodynamic description requires integrating over the infinite past $(-\infty,t)$.
Recently, significant strides have been made in understanding the structure of these generalized operators. A major breakthrough was achieved by Al-Refai and Fernandez [10], who provided a unified algebraic framework for Sonine operators via the method of conjugations. Their work elegantly demonstrates that many weighted operators on the half-line $0,\infty)$ can be constructed by conjugating standard derivatives with invertible linear operators, effectively unifying a wide range of fractional definitions under a common algebraic umbrella. However, their approach is primarily tailored to initial value problems using the Laplace transform on semi-axes. Consequently, the spectral analysis of “full history” operators on the unbounded domain $\mathbb{R}$—particularly those combining general Sonine kernels with time-dependent weights via the Fourier transform—remains an open problem. While abstract approaches to evolution equations have recently appeared (e.g., Lizama and Murillo-Arcila, 2025 [24]), an explicit harmonic analysis framework allowing for concrete calculations in these heterogeneous settings is still lacking.
This gap has led to a fragmentation of the field into a “zoo” of disconnected operators. Researchers must often choose between the power-law memory of Riemann-Liouville, the logarithmic scaling of Hadamard, the non-singular kernels of Caputo-Fabrizio, or tempered behaviors. While the Generalized Fractional Calculus provides the algebraic structure for arbitrary kernels, the physical modeling of aging viscoelasticity, inertial oscillation, and ultra-slow diffusion demands a framework that simultaneously handles general memory geometry and time-domain heterogeneity.
@@ 3.1 Admissible Kernels and Operators
Remark 3.2 (Diffusive vs. Oscillatory Regimes).
It is crucial to distinguish between two physical regimes covered by Definition 3.1:
  1. 1.
    The Diffusive Regime: Characterized by kernels $k(t)$ whose Laplace transforms are Complete Bernstein Functions (CBF). These operators generate semi-groups that preserve positivity, consistent with thermodynamic relaxation.
  2. 2.
    The Oscillatory Regime: Characterized by kernels (e.g., Bessel functions $J_{0}(t)$ associated with wave equations) that satisfy the general Sonine condition but do not belong to the CBF class. In this regime, the sign-preservation property is naturally lost, reflecting the inertial nature of the solution.
The Spectral Mapping Theorems presented in this work rely on the algebraic convolution structure and hold for both regimes, whereas the maximum principles are restricted to the diffusive case.
@@
  • Case $n>1$ (High-Order): This allows for kernels that are continuous at the origin (regular kernels). In this setting, the operator $\mathfrak{D}_{\psi,\omega}^{(\kappa)}$ acts as a weighted derivative of order $n$, involving the $n$-th derivative of the convolution.
Figure 1: Admissible Memory Kernels. Comparison of decay rates for different Sonine kernels. While classical fractional calculus relies on pure power-laws (blue), our framework incorporates tempered kernels (orange) and exponential decays (green), allowing for finite propagation speeds and thermodynamic consistency.
3.2 Weighted Weyl-Sonine Operators.
@@ 1. Explicit Form: The operator is defined pointwise by the weighted integral:
The construction in Definitions 3.7-3.8 hereditarily preserves the convolution algebra of the standard Sonine calculus, albeit non-trivially deformed by the diffeomorphism $\psi$ and the weight $\omega$.
@@
Table 1: Classification of special cases of the Weighted Weyl-Sonine Operator. The first block lists kernels from the Diffusive Regime (Complete Bernstein Functions, monotone dynamics), while the last entry illustrates the Oscillatory/Inertial Regime (non-Bernstein, wave-like dynamics).
@@ I. Diffusive Regime (Monotone / Bernstein)
II. Oscillatory Regime (Inertial / Non-Bernstein)
Bessel-Klein-Gordon $J_{0}(2\sqrt{\lambda z})$ $t$ 11
@@ 3.3 Spectral Analysis
Proposition 3.14 (Existence and Analytic Structure of Admissible Distributional Pairs).
Let $\Phi(s)$ be a function analytic in the right half-plane. A unique pair of tempered distributions $(k,\kappa)\in\mathcal{S}^{\prime}_{+}(\mathbb{R})\times\mathcal{S}^{\prime}_{+}(\mathbb{R})$ [^1] The notation $\mathcal{S}^{\prime}_{+}(\mathbb{R})$ denotes the subspace of tempered distributions supported on $0,\infty)$. This condition imposes physical causality (the memory kernel vanishes for $t<0$) and ensures the existence of the Laplace transform $\Phi(s)$. satisfying the generalized Sonine relation exists if and only if they are determined by the inverse Laplace transforms of (26):
@@
Moreover, regarding the physical admissibility within the distributional framework, we distinguish two levels of regularity based on the criteria established by Kochubei [17] and Hanyga [23]:
@@
    Singularity Condition (General Non-Locality): To generate a valid non-local derivative, the distribution $\kappa$ must not be regular at the origin. This is satisfied if $\Phi(s)$ behaves asymptotically as a Stieltjes function (e.g., $\Phi(s)\sim s^{\alpha-n}$), implying that $\kappa$ is a singular distribution.
@@
    Thermodynamic Consistency (Diffusive Subclass): Specifically for dissipative/diffusive systems, we require that the effective operator symbol $\Psi(s)=s^{n}\Phi(s)$ belongs to the class of Complete Bernstein Functions (CBF). This guarantees that the corresponding operator generates a positivity-preserving semi-group (see Remark 3.2).
  3. 3.
    Conjugate Regularity: If condition (2) holds (diffusive case with $n=1$), the conjugate kernel $k$ is also a Stieltjes distribution (typically a Locally Integrable Completely Monotone function), satisfying the distributional Sonine condition $k*\kappa=\delta$ within the class of singular measures.
@@
Algebraic Existence in $\mathcal{S}^{\prime}$. The generalized condition $(k*\kappa)(t)=t_{+}^{n-1}/(n-1)!$ corresponds in the Laplace domain to the algebraic equation $\hat{k}(s)\hat{\kappa}(s)=s^{-n}$. Since the Laplace transform is an isomorphism on the space of supported distributions, identifying $\hat{\kappa}(s)=\Phi(s)$ uniquely determines $\hat{k}(s)=(s^{n}\Phi(s))^{-1}$ as a tempered distribution.
Analytic Properties (Diffusive Case). Assume strictly that the operator governs a diffusive process, so $\Psi(s)=s^{n}\Phi(s)$ is a Complete Bernstein Function (CBF). By the properties of CBFs, its reciprocal $1/\Psi(s)$ is a Stieltjes function. The inverse Laplace transform of a Stieltjes function is known to be a Completely Monotone (CM) function (or measure). Thus, $k(t)=\mathcal{L}^{-1}\{1/\Psi(s)\}$ is a non-negative distribution with CM density. By duality, if $\Phi(s)$ is Stieltjes, then $\kappa$ is also a distribution with CM density (singular at the origin). This structural duality ensures that the resulting Weyl-Sonine operators preserve the cone of non-negative functions in $\mathcal{S}_{\psi,\omega}$, a necessary condition for the well-posedness of diffusion models. For oscillatory cases (where CBF conditions are relaxed), existence is guaranteed by the algebraic part, but positivity is not preserved. ∎
Remark 3.15 (Comparison with Classical Frameworks: Riemann-Liouville vs. Weyl).
It is important to contrast our construction with the classical theory developed by Luchko [2]. The classical approach primarily addresses Riemann-Liouville type operators defined on finite intervals $[0,T]$ acting on spaces of continuous functions (e.g., $C_{-1}$). In that setting, kernel pairs are strictly categorized by their integrability at the origin (classes $\mathcal{K}$ vs. $\mathcal{S}_{-1}$).
In contrast, our work generalizes these concepts to Weyl-type operators defined on the entire real line $\mathbb{R}$. This extension is non-trivial as it requires controlling the asymptotic behavior of functions at infinity. By constructing the Weighted Weyl-Sonine operators over the distribution space $\mathcal{S}^{\prime}_{\psi,\omega}$, we achieve two advancements:
  1. 1.
    We bypass the ad-hoc integrability conditions at the origin, as $\mathcal{S}^{\prime}$ naturally accommodates hypersingular functionals.
  2. 2.
    We extend the algebraic structure of the Sonine pairs to the full timeline, allowing for the spectral analysis of "aging" systems with infinite memory history, a feature not captured by the finite-start Riemann-Liouville definitions.
Consequently, our Proposition 3.14 adapts the physical admissibility criteria of Kochubei (Bernstein/Stieltjes) to the broader context of Weyl operators on tempered distributions.
We now proceed to the main spectral mapping theorem. It is worth noting that the classical fractional calculus algebra—characterized by the power law $s^{\alpha}$—is recovered here as a particular instance, now generalized by the spectral symbol $\Phi(s)$.
@@
Let $n\in\mathbb{N}$ and let $u\in\mathcal{S}_{\psi,\omega}(\mathbb{R})$. Let $(k,\kappa)$ be an admissible distributional Sonine pair with associated Laplace symbol $\Phi(s):=\hat{\kappa}(s)$. The Weighted Weyl-Sonine operators satisfy the following spectral relations:
@@
Since $v\in\mathcal{S}(\mathbb{R})$ and $k\in\mathcal{S}^{\prime}_{+}$, we apply the Convolution Theorem for tempered distributions. Using the Sonine condition in the frequency domain, $\hat{k}(i\xi)\hat{\kappa}(i\xi)=(i\xi)^{-n}$, we relate the symbol of $k$ to $\Phi$:
@@
The Generalized Spectral Mapping Theorem implies that the natural energy spaces for these operators are the Weighted Sonine-Sobolev spaces, defined via the magnitude of the full operator symbol $\Psi(i\xi)=(i\xi)^{n}\Phi(i\xi)$:
$\mathcal{H}_{\psi,\omega}^{\Phi}(\mathbb{R}):=\left\{u\in L^{2}_{\psi,\omega}(\mathbb{R}):\int_{\mathbb{R}}\left(1+|(i\xi)^{n}\Phi(i\xi)|^{2}\right)|\mathcal{F}_{\psi,\omega}u(\xi)|^{2}\,d\xi<\infty\right\}.$
@@ 3.4 The Generalized Marchaud Representation
Remark 3.21 (Abstract Operator Unification).
Before proceeding to the applications, it is worth noting that the structure presented here is an instance of a broader algebraic construction. If $\mathcal{T}:\mathcal{H}\to L^{2}(\mathbb{R})$ is any generic invertible linear operator, one can define an abstract Weyl-Sonine operator as $\mathfrak{D}_{\mathcal{T}}:=\mathcal{T}^{-1}\mathbb{D}^{(\kappa)}\mathcal{T}$. In this scenario, our "Weighted Fourier Transform" generalizes naturally to a Conjugated Fourier Transform defined by $\mathcal{F}_{\mathcal{T}}:=\mathcal{F}\circ\mathcal{T}$. Remarkably, the Spectral Mapping Theorem proved earlier holds invariantly:
$\mathcal{F}_{\mathcal{T}}(\mathfrak{D}_{\mathcal{T}}u)(\xi)=\Phi(i\xi)\mathcal{F}_{\mathcal{T}}u(\xi).$ (31)
Our choice of the specific weighted composition operator $\mathcal{T}u(t)=\omega(t)u(\psi(t))$ is motivated by its direct physical interpretation in modeling aging media (time-stretching) and tempered dynamics, providing a concrete realization of this abstract spectral equivalence.
4 Applications to Anomalous Relaxation, Diffusion, and Wave Dynamics
In this final section, we demonstrate the versatility of the Weighted Weyl-Sonine framework by solving evolution problems that arise in complex media. We address three distinct physical regimes: generalized relaxation (diffusive), distributed-order evolution (ultra-slow transitions), and inertial propagation (oscillatory dynamics).
@@ 4 Applications to Anomalous Relaxation, Diffusion, and Wave Dynamics
The theoretical framework developed in Section 3 supports operators of arbitrary order $n$. In this section, we primarily analyze the linear evolution equation $\mathfrak{D}_{\psi,\omega}^{(\kappa)}u+\lambda u=f$. This formulation is general enough to unify the description of purely dissipative systems (relaxation/diffusion) and systems with inertial energy exchange (waves), depending on the spectral choice of the kernel $\kappa$.
4.1 Spectral Solvability of Generalized Evolution Models
We consider the fundamental evolution equation governing the dynamics of a field $u\in\mathcal{S}_{\psi,\omega}$ (e.g., stress in a viscoelastic material, density in a porous medium, or amplitude of a wave packet) under the action of the generalized derivative.
@@ 4.1 Spectral Solvability of Generalized Evolution Models
where $\lambda>0$ is a structural parameter and $f\in L^{2}_{\psi,\omega}(\mathbb{R})$ is a source term. The physical nature of the solution is entirely determined by the spectral class of the kernel $\kappa$:
  • Diffusive Case: If $\kappa$ is a Bernstein function (Table 1, Block I), the system exhibits monotonic relaxation.
  • Oscillatory Case: If $\kappa$ allows for inertial effects (e.g., Bessel kernel, Table 1, Block II), the system supports wave propagation.
@@
Let $\Phi(s)=\mathcal{L}\{\kappa\}(s)$ be the spectral symbol of the derivative kernel. Assume the following Ellipticity Condition holds on the imaginary axis:
@@
Applying the Weighted Fourier Transform $\mathcal{F}_{\psi,\omega}$ to Eq. (32) and utilizing Theorem 3.16 (with $n=1$), the differential equation converts into an algebraic one in the frequency domain. Recall that for $n=1$, the symbol of the operator is $i\xi\Phi(i\xi)$:
@@
Under condition (33), the multiplier $m(\xi)=(i\xi\Phi(i\xi)+\lambda)^{-1}$ belongs to $L^{\infty}(\mathbb{R})$. Since the Weighted Fourier Transform is a unitary isomorphism between $L^{2}_{\psi,\omega}$ and $L^{2}$, the solution is uniquely determined in the spectral domain. The explicit time-domain representation (34) follows directly from the inverse structure $\mathcal{F}_{\psi,\omega}^{-1}=M_{\omega}^{-1}\circ Q_{\psi}\circ\mathcal{F}^{-1}$. ∎